# LOSC: LiDAR Open-voc Segmentation Consolidator

### v1

@@ Abstract
We study the use of image-based Vision-Language Models (VLMs) for open-vocabulary segmentation of lidar scans in driving settings. Classically, image semantics can be back-projected onto 3D point clouds. Yet, resulting point labels are noisy and sparse. We consolidate these labels to enforce both spatio-temporal consistency and robustness to image-level augmentations. We then train a 3D network based on these refined labels. This simple method, called LOSC, outperforms the SOTA of zero-shot open-vocabulary semantic and panoptic segmentation on both nuScenes and SemanticKITTI, with significant margins.
Keywords: LiDAR segmentation, LiDAR semantic segmentation, LiDAR panoptic segmentation, 3D scene understanding, VLM, open-vocabulary, zero-shot
002020202040404040nuScenes                      SemanticKITTIOV3D [19]GGSD [16]AFOV [17]LOSC (ours)-3.8-3.2-1.449.3SAL [22]LOSC (ours)35.2-6.5
@@
002020202040404040nuScenes              SemanticKITTISAL [22]LOSC (ours)-10.048.4SAL [22]SAL-4D [24]LOSC (ours)-7.6-1.632.4
@@
1 Introduction
First, some VLMs are trained with the sole supervision of image-text pairs, which makes dense tasks such as segmentation harder to learn [6]. Second, as training a VLM requires a huge quantity of data, some models are trained on data scrapped from the Web, which are thus subject to a distribution bias: they contain comparatively few samples from particular domains, e.g., medical images [7, 8]. This concerns both the visual contents (images) and the associated vocabulary (text). Third, “vision” in “VLM” generally only implies images, as 3D data (depth maps, point clouds, etc.) aligned with text is in much shorter supply than 2D data [9].
Therefore, most existing 3D VLMs (relating 3D data and text) actually operate via intermediate 2D data, leveraging a 2D VLM already trained on a large quantity of images: the 2D VLM is used to extract 2D information (labels or features), which is then back-projected onto 3D [9]. The 3D labels (or features) can then be directly employed as such, or used as pseudo-labels to train a 3D model, as done for the feature-based distillation of a 2D VFM into a 3D VFM [10].
@@ 1 Introduction
However, driving scenes suffers from the data shortage mentioned above: for this specific domain, there is relatively little visual content available associated with text [7, 8], and even less involving 3D data [9]. As a result, current VLMs still lag behind full supervision for the semantic and panoptic segmentation of driving scene images, and even more so for lidar scans [5, 11, 12].
Concretely, in driving scenes, VLM-based image segments are often missing or wrongly labeled, while object boundaries are imprecise. This label inaccuracy transfers to 3D point after back-projection. Additionally, some 3D points may get wrong 2D information due to parallax between camera and lidar sensors, and 3D points invisible from cameras (not in their field of view) do not get any information.
@@ 2 Related work
OpenScene [22] and CLIP2Scene [23] do so by leveraging point-pixel correspondences and training a 3D network to produces point features which aligns with the corresponding pixel representations extracted from a CLIP-like model. The process was then improved by exploiting, e.g., additional geometric regularization [24, 25], improving text queries [26, 27], or exploiting models such as SAM [20] for denoising [28, 29].
SAL [30] leverages SAM [20], to get instance masks from images, and MaskCLIP [31], to get per-mask CLIP features. These masks and features are lifted in 3D (thanks to point-pixel correspondences) and a 3D network is trained to predict these pairs of masks and features. SAL therefore enables zero-shot lidar panoptic segmentation. SAL was recently improved by operating on sequences rather than single scans in [32]. In this work, we train our 3D network for semantic segmentation and leverage ALPINE [33] to obtain panoptic labels.
@@ 3.1 Initial Zero-shot Semantic Segmentation of LiDAR data with the VLM
(In the following, we drop subscripts, superscripts and parameters if they are clear from the context.) These initial labels are both noisy (incorrect VLM labeling, camera-lidar parallax inconsistencies) and sparse (missing 2D VLM labeling, 3D points invisible from cameras) in the sense that a point $p\in P$ is given a label $L_{\text{vlm}}(p)\in\mathcal{C}\cup\{\emph{ignore}\}$, where ignore represents the lack of semantics. For a dataset like SemanticKITTI [35], this sparsity is particularly severe (cf. Section 4), because the 360 lidar is associated to only one camera, with a 90 field of view. In the following, we introduce three refinement strategies to improve both label quality and quantity.
@@ 3.2 Label Refinement with Time-Based Consolidation (TBC)
Due to resolution limit, faraway objects are hard to identify because they are perceived with only few pixels and points. However, as the ego vehicle drives and comes closer to them, their semantics becomes more apparent. To leverage semantic consistency over time, we aggregate each sequence $\mathcal{P}$ of labeled point clouds into the same coordinate system and enforce label consistency at voxel level by majority voting. Then, for any individual scan $P\in\mathcal{P}$, we assign to each point $p\in P$ the label of the voxel $p$ falls into. It produces scans $L_{\text{tim}}^{\mathcal{P}}=\text{TBC}(L_{\text{vlm}}^{\mathcal{P}})$ with time-consistent labels.
@@ 3.3 Label Refinement with Augmentation-Based Consolidation (ABC)
$L_{\text{vlm}}^{A}$ $=\cup_{I\in\mathcal{I}}\;\Pi_{I}(\text{VLM}(A(I),T_{\mathcal{C}}))$ (2)
$L_{\text{abc}}(p)$ $=\text{ABC}(L_{\text{vlm}}^{\mathcal{A}})(p)=\begin{cases}c&\text{if~{}~{}}\forall A\in\mathcal{A},L_{\text{vlm}}^{A}(p)=c\\ \emph{ignore}&\text{otherwise}\end{cases}$ (3)
@@ 3.4 Label Refinement Combining Augmentation- and Time-based Consolidation (ATC)
Even more than time-based consolidation, augmentation-based consolidation tends to increase label quality but also to reduce label quantity (cf. Section 4). To make sure points of each class remain both numerous enough and balanced enough, we introduce a way to combine $L_{\text{aug}}$ and $L_{\text{tim}}$.
@@
$\mathcal{C}_{\text{robust-aug}}$ =$=\{c\in\mathcal{C}\mid N_{\text{aug}}^{c}\geq N\text{~{}and~{}}N_{\text{aug}}^{c}/N_{\text{tim}}^{c}\geq\tau\}$ (4)
@@
[…]
@@ 4 Experiments
Method Segmentation Prompt Label coveragemAccmIoU%
Semantic Panoptic Minimal Rich
Grounded SAM [36] 68.0 32.8 21.1
Grounded SAM 2 [36] 75.3 46.6 24.3
@@
Table 1: Semantic segmentation benchmark of 2D open-vocabulary VLMs on the nuScenes validation set. Performance is evaluated using simple or rich prompts, with semantic or (merged) panoptic segmentation results where applicable.
@@
On the image side, for augmentation-based consolidation, we use the Albumentations library [37], applying 10 distinct image transformations: horizontal-flip, hue-saturation, blur, color-jitter, auto-contrast, sharpen, chromatic-aberration, emboss, fancy-pca, clahe. The appendix details the rationale of this choice. This set of augmentations is consistent in the sense that although a single disagreement is enough to discard the label of a point (cf. Eq. 3), empirical results show the robustness of pseudo-labeling (cf. appendix), leading to SOTA results (cf. Table 5).
Regarding 3D model pretraining, we leverage a self-supervised backbone to benefit from class-agnostic representations for our open-set segmentation approach. For our 3D network, we thus chose WaffleIron [38] (as did LeAP [34]), more precisely WI-48-768, for which self-supervised pretrained weights with ScaLR [10] are publicly available. This pretrained model is currently among SOTA self-supervised 3D backbones for lidars. When finetuning this network, we use the protocol of [10] but reduce the number of epochs to 10. For trainings without ScaLR initialization, we train the model from scratch as described in [38].
@@
Labeling consistency consistency pretraining mIoUmIoU%
@@
Third, our augmentation-based consolidation (ABC), directly applied on $L_{\text{vlm}}$ to produce $L_{\text{abc}}$, brings little gain on SemanticKITTI over just training from $L_{\text{vlm}}$, and even slightly degrades performance on nuScenes. Besides, refining $L_{\text{abc}}$
with TBC to produce $L_{\text{aug}}$ leads to similar results than with just $L_{\text{tim}}$. But ABC is actually not meant to operate on it own. The benefit of augmentation-based consolidation appears in fact when combining $L_{\text{aug}}$ and $L_{\text{tim}}$, where we obtain an extra boost of performance of 1.6 and 2.1 mIoU points on nuScenes and SemanticKITTI, respectively, compared to using $L_{\text{tim}}$ alone.
Iterative finetuning.
The results in Table 3 show that the quality of labels keeps improving after each round of finetuning, whether starting from $L_{\text{tim}}$ labels or $L_{\text{aug}{\otimes}\text{tim}}$. The best results are obtained when starting from $L_{\text{aug}{\otimes}\text{tim}}$, which shows that our iterative finetuning strategy is able to maintain the initial advantage provided by our most-refined labels. While a single iteration may provide a gain up to +2.4 mIoU pts (first iteration on SemanticKITTI), further iterating provides diminishing returns. What thus makes sense practically is to only iterate a few times. In LOSC, we perform 3 finetuning iterations.
@@
$L_{\text{tim}}=\text{TBC}(L_{\text{vlm}})$ 0 74.1 39.9 6.7 26.3
@@
Table 4: Quantity and quality of training labels computed on nuScenes and SemanticKITTI training sets.
@@
Comparison to state-of-the-art zero-shot semantic segmentation.
@@
LOSC outperforms the SOTA by +3.2 mIoU pts on nuScenes. It even outperforms methods that additionally use images at inference time by +1.4 mIoU point. The gain over the SOTA on SemanticKITTI is +6.5 mIoU points (but only a couple of other methods evaluate on this dataset). Note that we could not compare against LeAP [34], which evaluates with a specific protocol that is not fully specified. More specifically, the class mapping from the original and classical 19 classes of SemanticKITTI to their 11 classes is not provided, and the code is not publicly available.
Compared to some other approaches, we obtain those SOTA results without resorting to any prompt engineering, VLM tweaking, intricate architecture, complex training, or the use of images for inference. Besides, we only have very few and robust parameters, which are the same for all datasets. For instance, we set $\tau$ to 1/3 but any value between 1/4 and 1/2 would have given equal or similar results for both nuScenes and SemanticKITTI. Beyond these bounds, performance would start dropping a bit.
@@
Table 5: Semantic segmentation results on nuScenes and SemanticKITTI validation sets. $\dagger$ indicates the panoptic mIoU, which, in the case of nuScenes, slightly differs ($<$ 1 pt in general) from the classical semantic segmentation mIoU otherwise reported here (see appendix).
Comparison to state-of-the-art zero-shot panoptic segmentation
nuScenes SemanticKITTI
[…]
SAL [30] 38.4 47.8 77.2 47.5 29.2 33.9 24.8 32.3 66.8 17.4 30.2 28.7
SAL-4D [32] - - - - - - 30.8 - 76.9 25.5 34.6 -
[…]
@@
5 Limitations
@@ 6 Conclusion
Our approach actually is not limited to 2D VLMs. It could actually be also applied to any strong 2D semantic segmenter trained with full-supervision, alleviating the need for 3D annotation.
Appendix
We supplement the main paper with the performance on super-classes (Appendix A), classwise results (Appendix B), information about the quantity and quality of labeled points for training (Appendix C), details about prompts used in experiments (Appendix D), a detailed study concerning the use of augmentations (Appendix E), and qualitative results (Appendix F).
Appendix A Results with super-classes
Some methods like SAL [30] and LeAP [34] evaluate on broader super-classes. For instance, SAL merges car, motorcycle, truck, etc., into a single ‘vehicle’ class. In Table 7, we recall results with the original fine-grained classes and also report our results adopting the same merging strategy as in SAL [30], applying class merging during inference to predict super-class labels. (As said in the experiment section of the paper, no comparison is possible with LeAP, which does not provides its mapping from 19 classes into 11.)

### v2

@@ Abstract
We study the use of image-based Vision-Language Models (VLMs) for open-vocabulary segmentation of lidar scans in driving settings. Classically, image semantics can be back-projected onto 3D point clouds. Yet, resulting point labels are noisy and sparse. We consolidate these labels to enforce both spatio-temporal consistency and robustness to image-level augmentations. We then train a 3D network based on these refined labels. This simple method, called LOSC, outperforms the SOTA of zero-shot open-vocabulary semantic and panoptic segmentation on both nuScenes and SemanticKITTI, with significant margins. Code is available at https://github.com/valeoai/LOSC.
1 Introduction
First, some VLMs are trained with the sole supervision of image-text pairs, that makes dense tasks as segmentation harder to learn [23]. Second, as training a VLM requires lots of data, some models are trained on data scrapped from the Web, which are thus subject to a distribution bias: they contain comparatively few samples from particular domains, e.g., medical images [26, 8]. It concerns both the visual contents (images) and the associated vocabulary (text). Third, “vision” in “VLM” generally only implies images, as 3D data (depth maps, point clouds, etc.) aligned with text is in much shorter supply than 2D data [18].
020204040nuScenes          SemanticKITTIOV3D [19]GGSD [16]AFOV [17]LOSC (ours)-3.8-3.2-1.449.3SAL [22]LOSC (ours)35.2-6.5
@@ 1 Introduction
020204040nuScenes        SemanticKITTISAL [22]LOSC (ours)-10.048.4SAL [22]SAL-4D [24]LOSC (ours)-7.6-1.632.4
@@
Therefore, most existing 3D VLMs (relating 3D data and text) actually operate via intermediate 2D data, leveraging a 2D VLM already trained on a large quantity of images: the 2D VLM is used to extract 2D information (labels or features), which is then back-projected onto 3D [18]. The 3D labels (or features) can then be directly employed as such, or used as pseudo-labels to train a 3D model, e.g., for the feature-based distillation of a 2D VFM into a 3D VFM [22].
@@
However, driving scenes suffers from the data shortage mentioned above: for this specific domain, there is relatively little visual content available associated with text [26, 8], and even less involving 3D data [18]. As a result, current VLMs still lag behind full supervision for the semantic and panoptic segmentation of driving scene images, and even more so for lidar scans [10, 1, 31]. Concretely, in driving scenes, VLM-based image segments are often missing or wrongly labeled, while object boundaries are imprecise. This label inaccuracy transfers to 3D point after back-projection. Additionally, some 3D points may get wrong 2D information due to parallax between camera and lidar sensors, and 3D points invisible from cameras (not in their field of view) do not get any information.
@@ 2 Related work
OpenScene [20] and CLIP2Scene [7] do so using point-pixel correspondences and training a 3D network to produce point features which align with the corresponding pixel representations extracted from a CLIP-like model. The process was then improved by exploiting, e.g., additional geometric regularization [28, 27], improving text queries [29, 14], or exploiting models such as SAM [15] for denoising [5, 39].
SAL [19] uses SAM [15] to get instance masks from images, and MaskCLIP [9] to get per-mask CLIP features. The masks and features are lifted in 3D with point-pixel correspondences, and a 3D network is trained to predict the pairs of masks and features. SAL therefore enables zero-shot lidar panoptic segmentation. SAL was recently improved by operating on sequences rather than single scans in [35]. In this work, we train our 3D network for semantic segmentation and leverage ALPINE [25] to obtain panoptic labels.
@@ 3.1 Initial Zero-shot Semantic Segmentation of LiDAR data with the VLM
(In the following, we drop subscripts, superscripts and parameters if they are clear from the context.) These initial labels are both noisy (incorrect VLM labeling, camera-lidar parallax inconsistencies) and sparse (missing 2D VLM labeling, 3D points invisible from cameras) in the sense that a point $p\in P$ is given a label $L_{\text{vlm}}(p)\in\mathcal{C}\cup\{\emph{ignore}\}$, where ignore represents the lack of semantics. For a dataset like SemanticKITTI [2], this sparsity is particularly severe (cf. Sec. 4), because the 360 lidar is associated to only one camera, with a 90 field of view. In the following, we introduce three refinement strategies to improve both label quality and quantity.
@@ 3.2 Label Refinement with Time-Based Consolidation (TBC)
Due to resolution, faraway objects are hard to identify as they are perceived with only few pixels and points. However, as the ego vehicle drives and comes closer to them, their semantics becomes more apparent. To leverage semantic consistency over time, we aggregate each sequence $\mathcal{P}$ of labeled point clouds into the same coordinate system and enforce label consistency at voxel level by majority voting. Then, for any individual scan $P\in\mathcal{P}$, we assign to each point $p\in P$ the label of the voxel $p$ falls into. It produces scans $L_{\text{tim}}^{\mathcal{P}}=\text{TBC}(L_{\text{vlm}}^{\mathcal{P}})$ with time-consistent labels.
@@ 3.3 Label Refinement with Augmentation-Based Consolidation (ABC)
$L_{\text{vlm}}^{A}$ $=\bigcup_{I\in\mathcal{I}}\;\Pi_{I}(\text{VLM}(A(I),T_{\mathcal{C}}))$ (2)
$L_{\text{abc}}(p)$ $=\text{ABC}(L_{\text{vlm}}^{\mathcal{A}})(p)$
$=\begin{cases}c,&\text{if }\forall A\in\mathcal{A},\;L_{\text{vlm}}^{A}(p)=c\\ \emph{ignore},&\text{otherwise}\end{cases}$ (3)
@@ 3.4 Label Refinement Combining Augmentation- and Time-based Consolidation (ATC)
Even more than time-based consolidation, augmentation-based consolidation tends to increase label quality but also to reduce label quantity (cf. Sec. 4). To make sure points of each class remain both numerous enough and balanced enough, we introduce a way to combine $L_{\text{aug}}$ and $L_{\text{tim}}$.
@@
$\mathcal{C}_{\text{robust-aug}}$ $=\{c\in\mathcal{C}\mid N_{\text{aug}}^{c}\geq N\text{~and~}N_{\text{aug}}^{c}/N_{\text{tim}}^{c}\geq\tau\}$ (4)
@@
$L_{\text{aug}{\otimes}\text{tim}}(p)$ $=\text{comb}(L_{\text{aug}},L_{\text{tim}})(p)$
$=\begin{cases}L_{\text{aug}}(p),&L_{\text{aug}}(p)\in\mathcal{C}_{\text{robust-aug}}\\ L_{\text{tim}}(p),&\text{otherwise}\end{cases}$ (5)
@@ 4 Experiments
Method Segmentation Prompt Label cov.% mAcc mIoU
Sem. Pan. Min. Rich % %
GSAM [24] 68.0 32.8 21.1
GSAM 2 [24] 75.3 46.6 24.3
@@
Table 1: Semantic segmentation benchmark of 2D open-vocabulary VLMs on the nuScenes validation set. Performance is evaluated using minimal or rich prompts, with semantic or (merged) panoptic segmentation results where applicable. Label cov. is label coverage.
@@
On the image side, for augmentation-based consolidation, we use the Albumentations library [3], applying 10 distinct image transformations: horizontal-flip, hue-saturation, blur, color-jitter, auto-contrast, sharpen, chromatic-aberration, emboss, fancy-pca, clahe. The appendix details the rationale of this choice. This set of augmentations is consistent in the sense that although a single disagreement is enough to discard the label of a point (cf. Eq. 3), empirical results show the robustness of pseudo-labeling (cf. appendix), leading to SOTA results (cf. Tab. 6).
Regarding 3D model pretraining, we leverage a self-supervised backbone to benefit from class-agnostic representations for our open-set segmentation approach. As 3D network, we thus chose WaffleIron [21] (as did LeAP [11]), more precisely WI-48-768, for which self-supervised pretrained weights with ScaLR [22] are publicly available. This pretrained model is currently among SOTA self-supervised 3D backbones for lidars. When finetuning this network, we use the protocol of [22] but reduce the number of epochs to 10. For trainings without ScaLR initialization, we train the model from scratch as described in [21].
@@
Labeling consistency consistency pretraining mIoU% mIoU%
@@
Third, our augmentation-based consolidation (ABC), directly applied on $L_{\text{vlm}}$ to produce $L_{\text{abc}}$, brings little gain on SemanticKITTI over just training from $L_{\text{vlm}}$, and even slightly degrades performance on nuScenes. Besides, refining $L_{\text{abc}}$ with TBC to produce $L_{\text{aug}}$ leads to similar results than with just $L_{\text{tim}}$. But ABC is actually not meant to operate on it own. The benefit of augmentation-based consolidation appears in fact when combining $L_{\text{aug}}$ and $L_{\text{tim}}$, where we obtain an extra boost of performance of 1.6 and 2.1 mIoU points on nuScenes and SemanticKITTI, respectively, compared to using $L_{\text{tim}}$ alone.
@@
$L_{\text{tim}}=\text{TBC}(L_{\text{vlm}})$ 0 74.1 39.9 06.7 26.3
@@
Table 4: Quantity and quality of training labels computed on nuScenes and SemanticKITTI training sets. Note that some lidar points are not visible from the cameras and that the VLM does not assign labels to all pixels. Moreover, while the 6 nuScenes cameras observe the scene at 360°, the single camera of SemanticKITTI only covers about 1/4 of the lidar field of view. This explains the label coverage.
Iterative finetuning.
The results in Table 3 show that the quality of labels keeps improving after each round of finetuning, whether starting from $L_{\text{tim}}$ labels or $L_{\text{aug}{\otimes}\text{tim}}$. The best results are obtained when starting from $L_{\text{aug}{\otimes}\text{tim}}$, which shows that our iterative finetuning strategy is able to maintain the initial advantage provided by our most-refined labels. While a single iteration may provide a gain up to +2.4 mIoU pts, further iterating provides diminishing returns. What thus makes sense practically is to only iterate a few times. In LOSC, we perform 3 finetuning iterations.
@@
Comparison to SOTA zero-shot semantic segmentation.
@@
LOSC outperforms the SOTA by +3.2 mIoU pts on nuScenes. It even outperforms methods that additionally use images at inference time by +1.4 mIoU point. The gain over the SOTA on SemanticKITTI is +6.5 mIoU points (but only a couple of other methods evaluate on this dataset).
nuScenes SemanticKITTI
[…]
SAL [19] 38.4 47.8 77.2 47.5 29.2 33.9 24.8 32.3 66.8 17.4 30.2 28.7
SAL-4D [35] - - - - - - 30.8 - 76.9 25.5 34.6 -
LOSC (ours) 48.4 58.1 71.1 46.7 51.3 49.8 32.4 41.4 51.8 36.1 29.7 35.2
Note that we could not compare to LeAP [11], that evaluates with a specific protocol that is not fully described. In particular, the class mapping from the original and classical 19 classes of SemanticKITTI to their 11 classes is ambiguous and not provided in [11], while the code is not available.
Compared to some other approaches, we obtain those SOTA results without resorting to any prompt engineering, VLM tweaking, intricate architecture, complex training, or use of images for inference. Besides, we only have very few and robust parameters, which are the same for all datasets. For instance, we set $\tau$ to 1/3 but any value between 1/4 and 1/2 would have given equal or similar results for both nuScenes and SemanticKITTI. Beyond these bounds, performance would start dropping a bit.
We present in Figure 3 qualitative results on the nuScenes and SemanticKITTI validation sets.
@@
Table 6: Semantic segmentation results on nuScenes and SemanticKITTI validation sets. $\dagger$ indicates the panoptic mIoU, which, in the case of nuScenes, slightly differs (<< 1 pt in general) from the classical semantic segmentation mIoU otherwise reported here (see appendix).
Comparison to SOTA zero-shot panoptic segmentation.
@@
Ground truth LOSC
nuScenes
Refer to caption Refer to caption
nuScenes
Refer to caption Refer to caption
SemanticKITTI
Refer to caption Refer to caption
SemanticKITTI
Refer to caption Refer to caption
Figure 3: Qualitative results of semantic segmentation from the validation sets of nuScenes and SemanticKITTI. The color code used to represent each class is provided in Supplementary Material. A typical error with LOSC on both datasets is a confusion between different types of flat surfaces such as road/driveable surface, sidewalk and terrain. In SemanticKITTI, trunks are also systematically included in vegetation rather than considered as a separate class. These observations are consistent with the quantitative class-wise results provided in Supplementary Material.
Moving objects.
@@ 5 Conclusion
Our approach is not bound to 2D VLMs. It could be also applied to any strong 2D semantic segmenter trained with full-supervision, alleviating the need for 3D annotation.
Acknowledgments.
We acknowledge EuroHPC Joint Undertaking for awarding the project ID EHPC-REG-2024R02-234 access to Karolina, Czech Republic. This work was granted access to the HPC resources of IDRIS under the allocations AD011014946R1 made by GENCI. This work was also supported by the European Union’s Horizon Europe research and innovation programme under grant agreement No 101214398 (ELLIOT).
Appendix A Appendix
We supplement the main paper with the performance on super-classes (Appendix B), classwise results (Appendix C), information about the quantity and quality of labeled points for training (Appendix D), details about prompts used in experiments (Appendix E), and a detailed study concerning the use of augmentations (Appendix F).
Appendix B Results with super-classes
Some methods like SAL [19] and LeAP [11] evaluate on broader super-classes. For instance, SAL merges car, motorcycle, truck, etc., into a single ‘vehicle’ class. In Tab. 7, we recall results with the original fine-grained classes and also report our results adopting the same merging strategy as in SAL [19], applying class merging during inference to predict super-class labels. (As said in the experiment section of the paper, no comparison is possible with LeAP, which does not provides its mapping from 19 classes into 11.)